OpenStudy (scarlettfarra2000):

Find the Value of the expression 6+2(2*2-1) ---------- 3*4-11 simplify the answer I got 12 --- -7.6

OpenStudy (photon336):

Well for the top portion I agree with you and here's why you would use PEMDAS You would first do what's in the ( ) first (2*2-1) next thing it to evaluate the exponent 2*2 = 4 4-1 = 3 Then you get this (3)(2)+6 6+6 = 12

OpenStudy (photon336):

now the bottom you have 3*4-11 3*4 = 12-11 = 1 \[\frac{ 12 }{ 1 } = 12 \]

OpenStudy (photon336):

@ScarlettFarra2000

OpenStudy (scarlettfarra2000):

I 'm here

OpenStudy (scarlettfarra2000):

My computer running a little slow

OpenStudy (scarlettfarra2000):

Thank you